OpenStudy (anonymous):

The table lists data regarding the average salaries of several professional athletes in the years 1991 and 2001. A)use the date points to find a linear function that fits the date. use the function to predict the average salary in 2005 and 2010. Year….Av erage Salary. 1991 $264,000 2001 $1,450,000 A linear function that fits the date is s(x)=___. (Let x=the number of years since 1990, and let s= the average salary years from 1990). The predicted average salary for 2005 is $___ (round to the nearest whole number) The predicted average salary for 2010 is $___(round to the nearest whole number).

OpenStudy (anonymous):

did you manage to do something yourself?

OpenStudy (anonymous):

yes, i got the first part, but then got stuck at figuring the salary for 2005 and 2010.

OpenStudy (anonymous):

what function you got for first part?

OpenStudy (anonymous):

i got answer to 2005, but i don't know if its right

OpenStudy (anonymous):

s(x)=264,000+118600

OpenStudy (anonymous):

you missing an x

OpenStudy (anonymous):

(10)

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

you missing x in your function

OpenStudy (anonymous):

is it suppose to be in front the 264,000? I am not sure

OpenStudy (anonymous):

it's not correct either way

OpenStudy (anonymous):

s(10)=264,000+118600

OpenStudy (anonymous):

i just realised that i was incorrect in previous question ><

OpenStudy (anonymous):

uh oh...what happened?

OpenStudy (anonymous):

let's go back to it first :D

OpenStudy (anonymous):

how did you get that answer for first par? (now i am pretty sure it's really incorrect :P)

OpenStudy (anonymous):

hang on a sec..need to find those notes too.

OpenStudy (anonymous):

so regarding what i posted previously, my answer for the first part I came up with is y=144,000

OpenStudy (anonymous):

then I need to figure how to get the average salary for 2005 and 2010. Thats where I am lost

OpenStudy (anonymous):

no it's not 144000

OpenStudy (anonymous):

awe..ok.

OpenStudy (anonymous):

firstly you ALWAYS need to have x

OpenStudy (anonymous):

if there's no x it's always equal to that number, it never changes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how did you get that? you were doing ok in previous question, so you should do this one too :P

OpenStudy (anonymous):

m=264,000-1,450,000/1991-2001=1,186,000/10 1,186,000/10=118,600 y=264,000+1,180,000 y=144,000

OpenStudy (phi):

you found the correct slope. now use it and one of the points to find the equation y-y1= m(x-x1)

OpenStudy (anonymous):

again you didn't write x :D

OpenStudy (phi):

and you should use x (year -1990) instead of year

OpenStudy (anonymous):

near slope you ALWAYS have x

OpenStudy (anonymous):

and also 264,000 is incorrect

OpenStudy (anonymous):

phi can you explain how to get that

OpenStudy (anonymous):

ok now i don't understand why 264,00 is wrong and not sure where the X should be

OpenStudy (anonymous):

x should be near slope

OpenStudy (anonymous):

m=264,000-1,450,000/1991-2001=1,186,000/10 this is slope

OpenStudy (anonymous):

ok, so i got that part right

OpenStudy (anonymous):

no, 264000 is incorrect

OpenStudy (anonymous):

ok, I;ll go watch it...

OpenStudy (anonymous):

\[s(x)=mx+b\] \[s(x)=118600x+b\] we have 1991 years and 264000$ but we start as 1990=0, so 1991 is 1 so we have point (1,264000) (x,y) and we put it into function \[264000=118600\cdot1+b\] and \[s(x)=145400\] so first part answer is \[s(x)=118600x+145400\] do you understand?

OpenStudy (anonymous):

yes, i get the idea...so how would i figure for 2005?

OpenStudy (anonymous):

since years start from 1990, you need 2005-1990=15 so x is 15 and put x inside that function

OpenStudy (anonymous):

1991-2005 and 145400-264.000?

OpenStudy (anonymous):

oh ok..

OpenStudy (anonymous):

so what you got?

OpenStudy (anonymous):

2,181,000?

OpenStudy (phi):

show us how you got that

OpenStudy (anonymous):

yeah show us :P

OpenStudy (anonymous):

145000/15

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry 145400/15

OpenStudy (phi):

if you type 145000/15 in the google search window you get 9 666.66667

OpenStudy (anonymous):

put 15 instead of x s(x)=... s(15)=...

OpenStudy (anonymous):

ok, don't confuse me with that phi, :), i am confused enough to figure this stuff out

OpenStudy (anonymous):

i did....

OpenStudy (anonymous):

phi is right :P

OpenStudy (anonymous):

so write down how you did

OpenStudy (anonymous):

i know...sorry

OpenStudy (anonymous):

s(x)=118600x+145400 s(15)=118600*15+145400

OpenStudy (anonymous):

you haven't done this

OpenStudy (anonymous):

i kinda had it except for i didn't use the 118600

OpenStudy (anonymous):

why you didn't use that? it's function and you need to use everything :D

OpenStudy (anonymous):

well, i was getting confused and looking at my previous notes.

OpenStudy (anonymous):

so what about year of 2010?

OpenStudy (anonymous):

can you do it yourself?

OpenStudy (phi):

did you get s(15)? Ready to try 2010?

OpenStudy (anonymous):

ok, so let me figure for 2010 and ill let you guys know what i come up with OK..

OpenStudy (anonymous):

s(20)=118600*20+145400=2,517,400

OpenStudy (anonymous):

yay it's correct

OpenStudy (anonymous):

:) oh gosh...

OpenStudy (anonymous):

thanks you two for all the help, i feel a little stupid haha...but i gotta learn somehow..and sorry phi for being rude...i was really getting frustrated and ready to give up...

OpenStudy (phi):

khan has a bunch of videos about lines. Search for equation of line. They are all pretty short.

OpenStudy (anonymous):

you can watch all video of khan because they are all good :P

OpenStudy (anonymous):

thanks, will do...

OpenStudy (anonymous):

cool, :) will check that out and save it to my favorites

OpenStudy (anonymous):

you can also do some exercises here aswell

OpenStudy (anonymous):

nice..gonna need to do LOTs of exercises. other wise, I am back on here.

OpenStudy (anonymous):

have a great night you two, thanks again!

OpenStudy (anonymous):

u2 ;)